A solution is prepared such that [Li+ ] = 3.0 * 10-3 M while [PO4 3- ] = 2.0 * 10-3 M. If the Ksp of Li3PO4 is 2.37 * 10-11, will a precipitate form in the prepared solution? Show your complete solution. 


1
Expert's answer
2021-03-15T09:08:15-0400

For precipitate to be present, Qip > Ksp

Qip = 3 ×10-3 × 2 ×10-3 = 6 × 10-6


6 × 10-6 > Ksp

Hence, a precipitate will be formed